OpenStudy (anonymous):

Which of the following statements is true about lines 3y + 2x - 9 = 0 and 12y + 8x - 5 = 0? they are parallel They are perpendicular They are parallel to the x-axis They are parallel to the y-axis

OpenStudy (muzzack):

lets start using technology to solve this question 3y + 2x - 9 = 0 12y + 8x - 5 = 0 we have to solve for y in both equations

OpenStudy (muzzack):

lets start with the 1st equation, wait a moment ill brb

OpenStudy (anonymous):

k @Muzzack

OpenStudy (muzzack):

im back

OpenStudy (muzzack):

lets put all the nonvariable into the other side ------ 3y+2x=9 ------ now all we have to do is put 2x to the other side by subtracting ------ 3y=-2x+9 ------ easy...well now we have to divide 3 both sides to isolate y ------ \(y=-\frac{2}{3}x+3\)

OpenStudy (anonymous):

Okay so just divide three from both sizes or thats what that is ^? @Muzzack